------------------------------------------------------------ Date: Thu Aug 30 08:04:01 2001, Shift: 00:00-08:00 Supervisor: J.I.Jordre, with: Log book#8 pages 39 - 42 Shift Summary: ------------------------------------------------------------ 1. Store since 03:02, preceeded by a quench link interlock. Started with ZDC coincidence rates of ~80/s. 2. Run 4499: started 03:06, ended 03:42 when found that BBR 10, 18-25 were off. BBR voltages mounted after end of run. Run 4500: ~376.6k trigger 1 events Run 4501: still going strong 3. TOF Cal ADC Top 3: wide/strange ADC distribution with resulting wide pedestals. (Plots in logbook.) Chip in FEE card on T1 with bad pedestals and hot behaviour when collecting data. (Plots in logbook.) Should be replaced at next access. Problems with updating RUN DB. Currently last run stored in DB is 4489. For TPC pedestal calculations use alternativ Tpcped script in "Backup solutions" folder on operator desktop. ------------------------------------------------------------
This archive was generated by hypermail 2b30 : Thu Aug 30 2001 - 08:04:22 EDT